The insulation resistance of a kilometer of the cable having a diameter of 2 cm and an insulation thickness of 2 cm is 600 ohms. If the thickness of the insulation is increased to 3 cm, find the insulation resistance of the cable.

1
Expert's answer
2021-12-10T16:10:43-0500

Explanations & Calculations


  • The resistance can be thought proportional to the length and inversely to the area of the conductor according to the equation "\small R = \rho \large\frac{L}{A}" .

"\qquad\qquad
\begin{aligned}
\small R &\propto\small \frac{L}{A}\\
\small 600 & =\small \frac{L}{\pi(4^2-2^2)}\cdots\cdots(1)\\
\small R&=\small \frac{L}{\pi(5^2-2^2)}\cdots\cdots(2)\\
\\
\small (2) \div(1),\\
\\
\small \frac{R}{600}&=\small \frac{(4^2-2^2)}{(5^2-2^2)}
\end{aligned}"


  • Initial total thickness is "\small 2+2=4\,cm"
  • Final total thickness is "\small 2+3 = 5\,cm"
  • The area of the resistive insulation is the difference between the (conductor+insulator) - (core).


  • Now you can calculate the new resistance "\small R."